# FedSSA: Semantic Similarity-based Aggregation for Efficient Model-Heterogeneous Personalized Federated Learning

### v1

@@ Abstract
[…]
@@ 1 Introduction
As societies become increasingly aware of the importance of data privacy protection, centrally collecting large-scale data for model training has become less viable, especially under privacy regulations (e.g., GDPR Kairouz and others (2021)). To enable decentralized data owners to collaboratively train effective machine learning models while protecting privacy, federated learning (FL) McMahan and others (2017) has been proposed. In a typical FL system, a central FL server coordinates multiple data owners (FL clients) for model training. In each communication round, the server broadcasts its current global model to the clients. Each client then treats the received global model as its local model, and trains it on local private data. The updated local model is uploaded to the server. The server weighted averages (aggregates) the received local models to obtain an updated global model. The above steps are repeated until the global model converges. In FL, only model parameters are communicated between the server and clients. Potentially sensitive local data are not exposed, thereby protecting privacy.
@@
This model homogeneous (structurally-same) mode of FL requires all clients to train models with the same structures. It is not well-suited for scenarios involving data heterogeneity Zhu and others (2021a) (i.e., clients with local data following non-independent and identical distributions (non-IID)) and system heterogeneity Jiang and others (2022); Itahara and et al (2020) (i.e., clients with diverse communication and computation capabilities). Therefore, training a personalized FL model adaptively based on each client’s local data distribution and system capability helps improve performances Tan and others (2022a). Furthermore, enterprise FL clients are often reluctant to share their model structures with others due to intellectual property concerns. Thus, the field of Model-Heterogeneous Personalized Federated Learning (MHPFL), which allows each client to train a structurally-different personalized local model with a structure tailored to actual needs, has emerged.
@@
Existing MHPFL methods are mainly designed based on knowledge distillation, mutual learning, and model mixup. Knowledge distillation-based methods Li and Wang (2019); Lin and others (2020); Huang and others (2022b); Itahara and others (2023); Sattler and others (2022); Makhija and others (2022); Chang and others (2021) often rely on the availability of a public dataset that is large enough and follows a similar distribution with local private data for fusing knowledge from heterogeneous local models. However, such a public dataset is difficult to be obtained in practice. Other knowledge distillation-based methods Jeong and others (2018); Ahn and others (2019, 2020); Tan and others (2022b) without requiring public datasets often incur high computational overhead on clients and face the risk of privacy leakage. Mutual learning-based methods Shen and others (2020); Wu and others (2022) allow each client to train a large heterogeneous model and a small homogeneous model in a mutual learning manner. Only the small homogeneous models are uploaded to the server for aggregation. Training two models simultaneously leads to high computational overhead for clients. Model mixup-based methods Collins and others (2021); Liang and others (2020); Jang and others (2022) divide each client’s local model into a heterogeneous feature extractor and a homogeneous classifier. The server aggregates the homogeneous local classifiers to update the global classifier and then broadcasts it to clients. Each client replaces its local classifier with the updated global classifier or constructs a regularization term based on the distillation loss between local and global classifiers. They often ignore the semantic information of the classifier parameters, resulting in limited accuracy improvements.
@@
To enable the privacy-preserving, communication- and computation-efficient training of high performance models through MHPFL, we propose the Federated Semantic Similarity Aggregation (FedSSA) approach. It splits each client’s local model into two parts: 1) a heterogeneous feature extractor, and 2) a homogeneous classification header. Local-to-global knowledge transfer is achieved by semantic similarity-based classification header parameter aggregation. Since the parameters of classification headers from different clients corresponding to the same class are semantically similar, each client only needs to upload classification header parameters corresponding to locally seen classes. The FL server aggregates parameters from different classification headers by class to update the global header. Global-to-local knowledge transfer is achieved through parameter stabilization. To ensure stable convergence, we devise an adaptive parameter stabilization strategy that fuses local historical header parameters and the latest global header parameters for locally seen classes to update the local header of each FL client.
@@
Owing to the above design, FedSSA has the following advantages: 1) aggregating semantically similar head parameters corresponding to the same class stabilizes the decision boundaries and boosts performance; 2) the adaptive parameter stabilization strategy alleviates parameter shaking in the initial training rounds, thereby speeding up convergence; 3) the server and clients only need to transmit classification header parameters corresponding to seen classes, which is communication-efficient compared to transmitting the entire model; 4) low computational overhead for FL clients as they only fuse partial parameters from the global and local headers before local training; and 5) since only partial parameters are transmitted between the FL server and FL clients, no new privacy risk is introduced.
@@ 2 Related Work
Our work is closely related to MHPFL with complete model heterogeneity. These methods support flexible model heterogeneity, and can be divided into three categories: 1) knowledge distillation, 2) mutual learning, and 3) model mixup.
Knowledge Distillation-based MHPFL. Most knowledge distillation-based MHPFL methods rely on a public dataset (e.g., FedMD Li and Wang (2019), FedDF Lin and others (2020), FCCL Huang and others (2022b), DS-FL Itahara and others (2023), CFD Sattler and others (2022), FedHeNN Makhija and others (2022), Cronus Chang and others (2021), FSFL Huang and others (2022a), FedAUX Sattler and others (2021), FedKT Li and others (2021), Fed-ET Cho and others (2022), FedKEMF Yu and others (2022), FedGEMS Cheng and others (2021), KT-pFL Zhang and others (2021)) and allow the server to aggregate the logits or representations of heterogeneous local models on a public dataset to integrate knowledge from different clients. However, the public dataset is not always available. In addition, models only perform well when the public dataset follows a similar distribution to local data. Knowledge distillation-based MHPFL methods not requiring a public dataset are starting to emerge. FedGen Zhu and others (2021b) and FedZKT Zhang and others (2022) introduce zero-shot knowledge distillation into FL to train a generator for generating local representations or public shared datasets. However, time-consuming iterative training for the generator incurs high computational costs. In FD Jeong and others (2018), HFD Ahn and others (2019, 2020), FedProto Tan and others (2022b), FedGKT He and others (2020), each client uploads the (class-average) logits or representations of local data samples to the server for aggregation, the aggregated global logits or representations for each class are sent to clients. Each client calculates the distillation loss between the global logits/representation and each local data sample logits/representation belonging to the same class for training the local model. This incurs high computational costs for clients. Besides, uploading the logits/representations might compromise privacy.
@@ 4.1 Semantic Similarity-Based Aggregation
Figure 2: Examples of decay functions for $\mu^{t}$.
@@ 4.2 Adaptive Parameter Stabilization
Inspired by historical learning Li and others (2023), during global-to-local knowledge transfer, we exploit historical local headers to stabilize and speed up model convergence. Specifically, in the $t$-th training round, client $k$ fuses the seen-class parameters from its historical local header $\theta_{k}^{t-1,s}$ ($s\in S_{k}$) and the latest global header $\theta^{t-1,s}$ ($s\in S_{k}$) to produce a new local header $\hat{\theta}_{k}^{t,s}$. The unseen-class header parameters are still the historical local header $\theta_{k}^{t-1,s^{\prime}}$ ($s^{\prime}\in(S\backslash S_{k})$). The fusion of the seen-class parameters from the two headers is achieved by the proposed adaptive parameter stabilization strategy.
@@
The coefficient $\mu^{t}\in[0,1]$ of the historical local header is determined by the above piecewise function. In round $t\leq T_{stable}$ (during which parameter stabilization is required with a fixed hyperparameter for all clients), $\mu^{t}$ is determined by a decay function (e.g., Figure 2) with initial value $\mu_{0}\in(0,1]$. $\mu^{t}$ decreases as round $t$ rises. When $t>T_{stable}$, the reliability of the global header has been enhanced. Thus, we set $\mu^{t}=0$ to directly replace historical local seen-class parameters with the latest global seen-class header parameters. This prevents stateless local header parameters from slowing down convergence.
@@
These three key steps of uploading seen-class header parameters by class, aggregation by class, and parameter fusion by class enhance the personalization and stabilize the classification boundary for each structurally heterogeneous local FL model. They are repeated until local models from all clients converge. Each client’s final local model will then be used for inference. FedSSA is detailed in Algorithm 1 (Appendix A).
@@ 4.3 Discussion
Privacy. Since clients only share seen-class header parameters and classes with the server, inferring the complete model and even local data inversely is currently hard to achieve. Hence, FedSSA is consideered privacy-preserving.
Communication Cost. In the uplink communication, clients upload the seen-class local header parameters and classes to the server. In the downlink communication, clients receive the seen-class global header parameters from the server. Each client’s seen classes only require to be uploaded to the server in the first round. In addition, only the headers of the complete models are communicated between clients and the server in each round of FL. Therefore, FedSSA incurs lower communication cost than transmitting the complete models as in the cases of FedAvg based FL approaches.
@@ 5 Analysis
To analyze the convergence of FedSSA, we first introduce some additional notations. Let $e\in\{0,1,...,E\}$ denote a local iteration. $(tE+e)$ denotes the $e$-th iteration in the FL training round $(t+1)$. At $(tE+0)$ (i.e., the beginning of the round $(t+1)$), clients fuse historical local headers (in round $t$) and the latest global header (in round $t$) to update local headers. $(tE+1)$ is the first iteration in round $(t+1)$. $(tE+E)$ is the last iteration in round $(t+1)$.
@@
Bounded Variance of Classification Headers. The variance between the seen-class parameters $\theta_{k}^{s}(s\in S_{k})$ of client $k$’s local header $\mathcal{H}(\theta_{k})$ and the same seen-class parameters $\theta^{s}=\mathbb{E}_{k^{\prime}\in\mathcal{K}_{s}}[\theta_{k^{\prime}}^{s}](s\in S_{k})$ of the global header $\mathcal{H}(\theta)$ is bounded:
@@
Gradient bounded: $\mathbb{E}_{s\in S_{k}}[\|\nabla\mathcal{L}(\theta^{s})-\nabla\mathcal{L}(\theta_{k}^{s})\|_{2}^{2}]\leq\delta^{2}$.
@@
The detailed proof can be found in Appendix C.
@@
The proof can be found in Appendix D. Then we can derive the non-convex convergence rate of FedSSA as follows:
@@
Therefore, under FedSSA, an arbitrary client’s local model converges at a non-convex convergence rate of $\epsilon\sim\mathcal{O}(\frac{1}{T})$. The detailed proof can be found in Appendix E.
@@ 6 Experimental Evaluation
In this section, we compare FedSSA against $7$ state-of-the-art MHPFL approaches on two real-world datasets under various experiment conditions.
@@ 6.1 Experiment Setup
Datasets and Models. We evaluate FedSSA and baselines on two image classification datasets: CIFAR-10 and CIFAR-100  [^1] https://www.cs.toronto.edu/%7Ekriz/cifar.html Krizhevsky and others (2009). They are manually divided into non-IID datasets following the method specified in Shamsian and others (2021). For CIFAR-10, we assign only data from 2 out of the 10 classes to each client (non-IID: 2/10). For CIFAR-100, we assign only data from 10 out of the 100 classes to each client (non-IID: 10/100). Then, each client’s local data are further divided into the training set, the evaluation set, and the testing set following the ratio of 8:1:1. This way, the testing set is stored locally by each client, which follows the same distribution as the local training set. As shown in Table 4 (Appendix B), each client trains a CNN model with output layer dimensions (i.e., the last fully-connected layer) of $10$ or $100$ on CIFAR-10 and CIFAR-100 datasets, respectively. The dimensions of the representation layer (i.e., the second last fully-connected layer) are set to be $500$. Therefore, each classification header consists of $10(100)\times 500$ parameters.
@@ 6.2.2 Model-Heterogeneous PFL
In this scenario, we evenly assign $5$ models with different structures (Table 4, Appendix B) to clients. In practice, the id of the model assigned to client $k$ is determined by (client id $k\%5$). For FML and FedKD, we regard the $5$ structurally different models as large heterogeneous models and consider the smallest ‘CNN-5’ model as the small homogeneous model.
Mean Accuracy. Table 2 shows that FedSSA consistently achieves the highest mean test accuracy under all settings, outperforming the best baseline by $3.62\%$. Figure 8 (Appendix B) shows how the mean test accuracy of FedSSA and the two top-performing baselines in each FL setting of Table 2 varies with communication round, it presents that FedSSA converges to higher accuracies faster than others.
@@
Figure 4: Trade-off among test accuracy, communication cost, and computational overhead. The size of each marker reflects the corresponding computation FLOPs (1e12).
Trade-off Among Accuracy, Communication and Computational costs. Figure 4 reflects the trade-offs achieved by the comparison algorithms among accuracy, communication costs, and computational costs. The marker closer to the bottom-right-hand corner (higher accuracy and lower communication cost) of the figure and with a smaller size (lower computational cost) achieves the best trade-offs. Under all settings of $(N,C)$ (markers with different types), FedSSA achieves higher accuracies, and lower or similar communication and computational costs compared to the best-performing baseline, FedProto, thus demonstrating its capability of achieving the most advantageous trade-off.
Personalization Analysis. We extract every sample representation from each FL client under FedSSA and FedProto, respectively. Then, we leverage the T-SNE van der Maaten and Hinton (2008) tool to reduce the dimensionality of the extracted representations from $500$ to $2$, and visualize the results. Since CIFAR-100 includes 100 classes of samples, we focus on visualizing the results on CIFAR-10 (non-IID: 2/10) in Figure 5. It displays that most clusters in FedSSA and FedProto consist of representations from a client’s two seen classes of samples, which indicates that each client’s local heterogeneous model has strong personalization capability. The two seen class representations within most clusters under FedSSA and FedProto satisfy “intra-class compactness and inter-class separation”, reflecting that every client can classify its seen classes well under both algorithms. Generally, FedSSA performs better than FedProto.
Figure 5: T-SNE representation visualization for FedProto and FedSSA on the CIFAR-10 (Non-IID: 2/10) dataset.

### v3

@@ Abstract
Federated learning (FL) is a privacy-preserving collaboratively machine learning paradigm. Traditional FL requires all data owners (a.k.a. FL clients) to train the same local model. This design is not well-suited for scenarios involving data and/or system heterogeneity. Model-Heterogeneous Personalized FL (MHPFL) has emerged to address this challenge. Existing MHPFL approaches often rely on a public dataset with the same nature as the learning task, or incur high computation and communication costs. To address these limitations, we propose the Federated Semantic Similarity Aggregation (FedSSA) approach for supervised classification tasks, which splits each client’s model into a heterogeneous (structure-different) feature extractor and a homogeneous (structure-same) classification header. It performs local-to-global knowledge transfer via semantic similarity-based header parameter aggregation. In addition, global-to-local knowledge transfer is achieved via an adaptive parameter stabilization strategy which fuses the seen-class parameters of historical local headers with that of the latest global header for each client. FedSSA does not rely on public datasets, while only requiring partial header parameter transmission to save costs. Theoretical analysis proves the convergence of FedSSA. Extensive experiments present that FedSSA achieves up to $3.62\%$ higher accuracy, $15.54$ times higher communication efficiency, and $15.52$ times higher computational efficiency compared to 7 state-of-the-art MHPFL baselines.
@@ 1 Introduction
As societies become increasingly aware of the importance of data privacy protection Yi and others (2020), centrally collecting large-scale data for model training has become less viable, especially under privacy regulations. To enable decentralized data owners to collaboratively train effective machine learning models while protecting privacy, federated learning (FL) McMahan and others (2017) has been proposed. In a typical FL system, a central FL server coordinates multiple data owners (FL clients) for model training. In each communication round, the server broadcasts its current global model to the clients. Each client then treats the received global model as its local model, and trains it on local private data. The updated local model is uploaded to the server. The server weighted averages (aggregates) the received local models to obtain an updated global model. The above steps are repeated until the global model converges. In FL, only model parameters are communicated between the server and clients. Potentially sensitive local data are not exposed, thereby protecting privacy Yi and others (2023b); Shi et al. (2023); Liu et al. (2024).
Figure 1: Feature extractor and classification header.
@@
This model homogeneous mode of FL requires all clients to train models with the same structures. It is not well-suited for scenarios involving data heterogeneity Zhu and others (2021a) (i.e., clients with local data following non-independent and identical distributions (non-IID)) and system heterogeneity Jiang and others (2022); Yi and others (2024a); Yu et al. (2011); Zhang and others (2023b, 2020, a, 2024) (i.e., clients with diverse communication and computation capabilities). Therefore, training a personalized FL model adaptively based on each client’s local data distribution and system capability helps improve performances Tan and others (2022a); Yi and others (2023d, 2024b). Furthermore, enterprise FL clients are often reluctant to share their model structures with others due to intellectual property concerns. Thus, the field of Model-Heterogeneous Personalized Federated Learning (MHPFL) Yi et al. (2023); Yi and others (2023c, a, 2024c), which allows each client to train a heterogeneous personalized local model with a structure tailored to actual needs, has emerged.
@@
Existing MHPFL methods are mainly designed based on knowledge distillation, mutual learning, and model mixup. Knowledge distillation-based methods Li and Wang (2019); Lin and others (2020); Huang and others (2022b); Itahara and others (2023); Sattler and others (2022); Makhija and others (2022); Chang and others (2021) often rely on the availability of a public dataset that is large enough and follows a similar distribution with local private data for fusing knowledge from heterogeneous local models. However, such a public dataset is difficult to obtain in practice. Other knowledge distillation-based methods Jeong and others (2018); Ahn and others (2019, 2020); Tan and others (2022b) without requiring public datasets often incur high computational overhead on clients and face the risk of privacy leakage. Mutual learning-based methods Shen and others (2023); Wu and others (2022) allow each client to train a large heterogeneous model and a small homogeneous model in a mutual learning manner. Only the small homogeneous models are uploaded to the server for aggregation. Training two models simultaneously leads to high computational overhead for clients. Model mixup-based methods Collins and others (2021); Liang and others (2020); Jang and others (2022) divide each client’s local model into a heterogeneous feature extractor and a homogeneous classifier. The server aggregates the homogeneous local classifiers to update the global classifier and then broadcasts it to clients. Each client replaces its local classifier with the updated global classifier or constructs a regularization term based on the distillation loss between local and global classifiers. They often ignore the semantic information of the classifier parameters, resulting in limited accuracy improvements.
@@
To enable the privacy-preserving, communication- and computation-efficient training of high-performance models through MHPFL, we propose the Federated Semantic Similarity Aggregation (FedSSA) approach for supervised classification tasks, which have been widely applied in fields like cancer detection for medical diagnosis and object recognition in autonomous driving Kairouz and others (2021). As depicted in Figure 1, it splits each client’s local model into two parts: 1) a heterogeneous feature extractor, and 2) a homogeneous classification header. Local-to-global knowledge transfer is achieved by semantic similarity-based classification header parameter aggregation. Since the parameters of classification headers from different clients corresponding to the same class are semantically similar, each client only needs to upload classification header parameters corresponding to locally seen classes. The FL server aggregates parameters from different classification headers by class to update the global header. Global-to-local knowledge transfer is achieved through parameter stabilization. To ensure stable convergence, we devise an adaptive parameter stabilization strategy that fuses local historical header parameters and the latest global header parameters for locally seen classes to update the local header of each FL client.
@@
Owing to the above design, FedSSA has the following advantages: 1) aggregating semantically similar head parameters corresponding to the same class stabilizes the decision boundaries and boosts performance; 2) the adaptive parameter stabilization strategy alleviates parameter shaking in the initial training rounds, thereby speeding up convergence; 3) the server and clients only need to transmit classification header parameters corresponding to seen classes, which is communication-efficient compared to transmitting the entire model; 4) low computational overhead for FL clients as they only fuse partial parameters from the global and local headers before local training; and 5) since only partial parameters are transmitted between the FL server and FL clients, no new privacy risk is introduced. In short, compared with existing methods, FedSSA improves model performance, and communication and computation efficiency simultaneously.
@@ 2 Related Work
Our work is closely related to MHPFL with complete model heterogeneity. These methods support flexible model heterogeneity, and are divided into the following three categories.
Knowledge Distillation-based MHPFL. Most knowledge distillation-based MHPFL methods rely on a public dataset (e.g., FedMD Li and Wang (2019), FedDF Lin and others (2020), FCCL Huang and others (2022b), DS-FL Itahara and others (2023), CFD Sattler and others (2022), FedHeNN Makhija and others (2022), Cronus Chang and others (2021), FSFL Huang and others (2022a), FedAUX Sattler and others (2021), FedKT Li and others (2021), Fed-ET Cho and others (2022), FedKEMF Yu and others (2022), FedGEMS Cheng and others (2021), KT-pFL Zhang and others (2021)) to fuse information of local heterogeneous models by knowledge distillation. However, the public dataset is not always available. To avoid relying public datasets, FedGen Zhu and others (2021b) and FedZKT Zhang and others (2022) train a generator for generating local representations or public shared datasets, which is time-consuming and computation-intensive. In FD Jeong and others (2018), HFD Ahn and others (2019, 2020), FedProto Tan and others (2022b), FedGKT He and others (2020), each client uploads the (class-average) logits or representations of local data to the server, the aggregated global logits or representations for each class are sent to clients for local knowledge distillation which incurs high computational costs for clients. Besides, uploading the logits/representations might compromise privacy.
@@ 4.1 Semantic Similarity-Based Aggregation
Figure 3: Decay functions for $\mu^{t}$. $\cos(\cdot)\in[0,\pi/2]$ is a smooth decay function. It leads to a stable decrease of $\mu$.
@@ 4.2 Adaptive Parameter Stabilization
Inspired by historical learning Li and others (2023) which tends to re-use historical knowledge from old models to avoid forgetting, during global-to-local knowledge transfer, we exploit historical local headers to stabilize and speed up model convergence. Specifically, in the $t$-th training round, client $k$ fuses the seen-class parameters from its historical local header $\theta_{k}^{t-1,s}$ ($s\in S_{k}$) and the latest global header $\theta^{t-1,s}$ ($s\in S_{k}$) to produce a new local header $\hat{\theta}_{k}^{t,s}$. The unseen-class header parameters are still the historical local header $\theta_{k}^{t-1,s^{\prime}}$ ($s^{\prime}\in(S\backslash S_{k})$). The fusion of the seen-class parameters from the two headers is achieved by the proposed adaptive parameter stabilization strategy.
@@
The coefficient $\mu^{t}\in[0,1]$ of the historical local header is determined by the above piecewise function. In round $t\leq T_{stable}$ (a fixed hyperparameter for all clients in FL), $\mu^{t}$ is determined by a decay function (e.g., Figure 3) with initial value $\mu_{0}\in(0,1]$. $\mu^{t}$ decreases as round $t$ rises. When $t>T_{stable}$, the reliability of the global header has been enhanced. Thus, we set $\mu^{t}=0$ to directly replace historical local seen-class parameters with the latest global seen-class header parameters. This prevents stateless local header parameters from slowing down convergence.
@@
These three key steps of uploading seen-class header parameters by class, aggregation by class, and parameter fusion by class enhance the personalization and stabilize the classification boundary for each heterogeneous local model. They are repeated until local models from all clients converge. Each client’s final local model is used for inference. FedSSA is detailed in Algorithm 1 (Appendix A).
@@ 4.3 Discussion
Privacy. When a client uploads its seen-class header parameters, it can maintain the seen-class header parameters and replace the unseen-class header parameters with 00, while still uploading the entire header to the server. The server aggregates non-zero local header parameters according to the corresponding classes (ordinates), and then broadcasts the updated global header to the clients. Since only the header parameters are transmitted, local data privacy can be preserved.
Communication Cost. As stated above, only the parameters of headers are communicated between clients and the server in each round of FL. Therefore, FedSSA incurs lower communication costs than transmitting the complete models as in the cases of FedAvg based FL approaches.
@@ 5 Analysis
To analyze the convergence of FedSSA, we first introduce some additional notations. Let $e\in\{0,1,...,E\}$ denote a local iteration. $(tE+e)$ denotes the $e$-th local iteration in the FL training round $(t+1)$. At $(tE+0)$ (i.e., the beginning of the round $(t+1)$), clients fuse historical local headers (in round $t$) and the latest global header (in round $t$) to update local headers. $(tE+1)$ is the first iteration in round $(t+1)$. $(tE+E)$ is the last iteration in round $(t+1)$.
@@
Bounded Variance of Classification Headers. The variance between the seen-class parameters $\theta_{k}^{s}(s\in S_{k})$ of client $k$’s local header $\mathcal{H}(\theta_{k})$ and the same seen-class parameters $\theta^{s}=\mathbb{E}_{k^{\prime}\in\mathcal{K}_{s}}[\theta_{k^{\prime}}^{s}](s\in S_{k})$ of the global header $\mathcal{H}(\theta)$ is bounded Yi et al. (2023):
@@
Gradient bounded: $\mathbb{E}_{s\in S_{k}}[\|\nabla\mathcal{L}(\theta^{s})-\nabla\mathcal{L}(\theta_{k}^{s})\|_{2}^{2}]\leq\delta_{1}^{2}$.
@@
The detailed proof can be found in Appendix B.
@@
The proof can be found in Appendix C. Then we can derive the non-convex convergence rate of FedSSA as follows:
@@
Therefore, under FedSSA, an arbitrary client’s local model converges at a non-convex convergence rate of $\epsilon\sim\mathcal{O}(\frac{1}{T})$. The detailed proof can be found in Appendix D.
@@ 6 Experimental Evaluation
In this section, we compare FedSSA  [^1] https://github.com/LipingYi/FedSSA against $7$ state-of-the-art MHPFL approaches on two real-world datasets under various experiment conditions. and baselines with Pytorch on four NVIDIA GeForce RTX 3090 GPUs with 24G memory.
@@ 6.1 Experiment Setup
Datasets and Models. We evaluate FedSSA and baselines on two image classification datasets: CIFAR-10 and CIFAR-100  [^2] https://www.cs.toronto.edu/%7Ekriz/cifar.html Krizhevsky and others (2009). They are manually divided into non-IID datasets following the method specified in Shamsian and others (2021). For CIFAR-10, we assign only data from 2 out of the 10 classes to each client (non-IID: 2/10). For CIFAR-100, we assign only data from 10 out of the 100 classes to each client (non-IID: 10/100). Then, each client’s local data are further divided into the training set, the evaluation set, and the testing set following the ratio of 8:1:1. This way, the testing set is stored locally by each client, which follows the same distribution as the local training set. As shown in Table 3 (Appendix E), each client trains a CNN model with output layer dimensions (i.e., the last fully-connected layer) of $10$ or $100$ on CIFAR-10 and CIFAR-100 datasets, respectively. The dimensions of the representation layer (i.e., the second last fully-connected layer) are set to be $500$. Therefore, each classification header consists of $10(100)\times 500$ parameters.
@@ 6.2.2 Model-Heterogeneous PFL
We evenly assign $5$ models (model IDs: $1-5$) with different structures (Table 3, Appendix E) to clients. In practice, the id of the model assigned to client $k$ is determined by (client ID $k\%5$). For FML and FedKD, we regard the $5$ structurally different models as large heterogeneous models and consider the smallest ‘CNN-5’ model as the small homogeneous model.
Mean Accuracy. Table 2 shows that FedSSA consistently achieves the highest mean test accuracy under all settings, outperforming the best baseline by $3.62\%$. Figure 10 (Appendix E) shows how the mean test accuracy of FedSSA and the two top-performing baselines in each FL setting of Table 2 varies with communication round, it presents that FedSSA converges to higher accuracies faster than others.
@@
Figure 5: Trade-off among test accuracy, communication cost, and computational overhead. The sizes of each marker (dots of varying sizes) reflect the corresponding computation FLOPs (1e12).
Trade-off Among Accuracy, Communication and Computational costs. Figure 5 reflects the trade-offs achieved by the comparison algorithms among accuracy, communication costs, and computational costs. The marker closer to the bottom-right-hand corner (higher accuracy and lower communication cost) of the figure and with a smaller size (lower computational cost) achieves the best trade-offs. Under all settings of (N, C) (markers with different types), FedSSA achieves higher accuracies, and lower or similar communication and computational costs compared to the best-performing baseline, FedProto, thus demonstrating its capability of achieving the most advantageous trade-off. Since we still upload the entire processed header to the server. One header has more parameters than the uploaded average representations of seen classes in FedProto, i.e., FedSSA has a higher per-round communication cost. Although it converges faster, its total communication costs are slightly higher than FedProto.
Personalization Analysis. We extract every sample representation from each FL client under FedSSA and FedProto, respectively. Then, we leverage the T-SNE van der Maaten and Hinton (2008) tool to reduce the dimensionality of the extracted representations from $500$ to $2$, and visualize the results. Since CIFAR-100 includes 100 classes of samples, we focus on visualizing the results on CIFAR-10 (non-IID: 2/10) in Figure 6. It displays that most clusters in FedSSA and FedProto consist of representations from a client’s two seen classes of samples, indicating that each client’s local model has strong personalization capability. Generally, FedSSA performs better than FedProto.
[…]